Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-08-29

The night's conditions were not photometric

Observers: Joe Lyman, David Homan
Support team: Lluís Galbany
TAT: Joe Anderson

Calibrations: EFOSC: bias, arcs

Conditions: Light clouds at the start, increasing during the night. Cloud coverage forced closure during the night. 

UT      Target          Type                    Comment
==============================================================================================
22:55   PS17egl              Follow up               K imaging; no detection in individual frames
00:20   -                           -                               Switch to EFOSC
00:40   PS17ehh            Classification         Object faint; red continuum apparent H_alpha emission at z~0.022
01:10   -                           -                               Repayment for exchanged time
01:50   Feige 110           Standard
02:15   AT2017gie         Classification         Target quite faint, spectrum very noisy (clouds), no clear ID
02:25   -                           -                               Image analysis
02:45   ATLAS17jym     Classification          Non detection (clouds)
03:00   ATLAS17jua      Classification          Noisy spectrum, possible Ia SN
03:30   -                          -                                Closed due to clouds
07:45   -                          -                                Reopened (switched to SOFI in the interim)
07:50   AT2017gax       Follow Up                HJK imaging
08:10   AT2017gax       Follow Up                SOFI BG
08:40   Hip025573       Telluric Standard
08:45   Feige 110         Flux Standard
08:55   AT2017fzw       Follow Up                SOFI BG
10:00   Hip020677       Telluric Standard
10:05   AT2017fzw        Follow Up               HJK imaging

Comments